# Yoshinaga’s criterion and the Varchenko–Gelfand algebra

### v1

@@
Abstract. A central question in the theory of hyperplane arrangements is when the complement of a complex arrangement is $K(\pi,1)$. Yoshinaga provides a necessary condition for complexifications of real arrangements, which he formulates purely combinatorially. We show that Yoshinaga’s criterion is equivalent to a natural statement about the Varchenko–Gelfand ring, which in practice allows for fast calculation. We conclude with an investigation of the relationships between various properties of arrangements, including Yoshinaga’s criterion and the $K(\pi,1)$ property.
@@ 1 Introduction
It is a long-standing open problem to determine which complex hyperplane arrangement complements are $K(\pi,1)$, meaning that their higher homotopy groups vanish. In the case where the hyperplane arrangement is the complexification of a real hyperplane arrangement, the homotopy type of the complement is determined by the associated oriented matroid [Sal87, Theorem 1], therefore the $K(\pi,1)$ problem must have a combinatorial answer. See [FR00, FR87a, Yos24b] for surveys and partial results.
Recently, Yoshinaga gave a necessary (but not sufficient) condition for a complexified arrangement complement to be $K(\pi,1)$ [Yos24a], which is formulated in terms of the oriented matroid. In the first part of this paper, we give an algebraic reformulation of Yoshinaga’s criterion, which we now describe.
Let $\mathcal{A}$ be a finite set of hyperplanes in a real vector space $V$, and let $\mathbb{F}$ be any field. The Varchenko–Gelfand algebra $\operatorname{VG}(\mathcal{A},\mathbb{F})$ is by definition the ring of locally constant $\mathbb{F}$-valued functions on the complement of the union of hyperplanes. This is a boring ring (it is isomorphic to a direct sum of one copy of $\mathbb{F}$ for each chamber), but it admits an interesting presentation whose generators are the Heaviside functions: there are two such functions for each hyperplane, taking the value 1 on one side of the hyperplane and 0 on the other side. The ring $\operatorname{VG}(\mathcal{A},\mathbb{F})$ is filtered, with the $p^{\text{th}}$ filtered piece consisting as functions that can be expressed as polynomials of degree at most $p$ in the Heaviside functions. The associated graded algebra, which is also called the Cordovil algebra, is isomorphic to the cohomology ring of the complement of the union of the subspaces $H\otimes{\mathbb{R}}^{3}\subset V\otimes{\mathbb{R}}^{3}$ [Mos17, DBPW24]. We say that the Varchenko–Gelfand algebra is quadratic if all relations among the Heaviside functions are generated by those of degree at most 2. Similarly, we say that the Cordovil algebra is quadratic if all relations among the corresponding generators are generated by those of degree 2. Our main results (Theorem 2.4 and Corollary 2.5) say that the following implications hold: $\operatorname{C}(\mathcal{A},\mathbb{F})$ is quadratic $\Longrightarrow$ $\operatorname{VG}(\mathcal{A},\mathbb{F})$ is quadratic $\Longleftrightarrow$ $\mathcal{A}$ satisfies Yoshinaga’s criterion.
@@
At first sight, Yoshinaga’s criterion (which is formulated combinatorially) might seem easier to work with than the condition that $\operatorname{VG}(\mathcal{A},\mathbb{F})$ is quadratic. In fact, our experience is that the algebraic condition is much faster to check, since computers are very good at using Gröbner bases to determine whether or not two ideals are equal. For example, let $\mathcal{A}$ be the arrangement whose normal vectors are given by the columns of the following matrix:
@@
It took about 1.65 seconds for Macaulay2 to determine that the Varchenko–Gelfand ideal is not quadratic. On the other hand, it took 3 days, 23 hours, 21 minutes, and 22 seconds for Sage to check Yoshinaga’s criterion directly. [^4] Our Sage implementation could admit many improvements; but, even with considerable effort, it is unlikely that we could beat the time of the easy Macaulay2 calculation.
Our primary motivation for Theorem 2.4 is to be able to perform fast calculations, and in particular to probe the question of how close Yoshinaga’s criterion is to the $K(\pi,1)$ property. The following examples illustrate the type of calculations that are made possible by our result.
@@
Let $\mathcal{A}$ be the arrangement in $\mathbb{R}^{6}$ consisting of all hyperplanes of the from $x_{i}=x_{j}$ for $1\leq i<j\leq 6$, together with the hyperplanes $x_{i}+x_{j}=0$ whenever jij-i is prime. (This is an intentionally unmotivated condition that is meant to produce a somewhat random arrangement lying in between the Coxeter arrangements of type $A_{5}$ and $D_{6}$.) The ring $\operatorname{VG}(\mathcal{A},{\mathbb{Q}})$ is not quadratic, and therefore $\mathcal{A}$ is not $K(\pi,1)$. This can be checked in Macaualy2 in about 30 seconds.
@@
When $t\in\{-1,0,1\}$, these arrangements have quadratic Varchenko–Gelfand algebras, and therefore satisfy Yoshinaga’s criterion. Edelman–Reiner, however, show that these arrangements are not $K(\pi,1)$ [ER95, Theorem 2.1].
@@
This is the smallest known non-tame arrangement (see [Abe25] for background). Then $\operatorname{VG}(\mathcal{A};{\mathbb{Q}})$ is quadratic, thus $\mathcal{A}$ satisfies Yoshinaga’s criterion. It is not known to the authors whether or not $\mathcal{A}$ is $K(\pi,1)$. Yoshinaga’s criterion gives supporting evidence that it could be.
Section 3 is devoted to relating Yoshinaga’s criterion to other algebraic, topological, and combinatorial conditions. We define what it means for a matroid to be chordal, generalizing the notion of a chordal graph. We then say that $\mathcal{A}$ is chordal if its underlying matroid is chordal. We prove that every real, chordal arrangement satisfies Yoshinaga’s criterion (Theorem 3.2). We also provide a proof (communicated to us by Paul Mücksch) that every arrangement satisfying Yoshinaga’s criterion is formal. The converses to these two theorems are false (Example 3.3 and Remark 3.8), but for graphical arrangements, chordality, formality, Yoshinaga’s criterion, and the $K(\pi,1)$ property are all equivalent (Corollary 3.7). Finally, we provide a chart that illustrates the implications between various properties known to be related to the $K(\pi,1)$ property, including chordality, formality, and more.
@@ Acknowledgements
The authors would like to thank Nick Addington, Mike Falk, and Paul Mücksch for their valuable contributions.
2 Yoshinaga’s criterion and the Varchenko–Gelfand algebra
@@
2.1 Yoshinaga’s criterion
We begin by choosing coorientations of each element of $\mathcal{A}$. That is, for each $H\in\mathcal{A}$, we write $H^{+}$ to denote one of the two connected components of $V\setminus H$, and $H^{-}$ to denote the other one. For any sign vector $\epsilon\in\{\pm\}^{\mathcal{A}}$ and any subset $S\subset\mathcal{A}$, let
@@ 2.1 Yoshinaga’s criterion
We say that $\mathcal{A}$ satisfies Yoshinaga’s criterion if $\sigma_{2}=\sigma_{r}$. This definition is motivated by the following result [Yos24a, Theorem 5.1(2)].
@@
If $M_{2}(\mathcal{A})$ is $K(\pi,1)$, then $\mathcal{A}$ satisfies Yoshinaga’s criterion.
@@ 2.2 The Varchenko–Gelfand algebra
In order to relate Yoshinaga’s criterion to the Varchenko–Gelfand ring, we introduce a family of smaller ideals that sit inside the kernel of $\varphi$. For any $k$, we define the $k^{\text{th}}$ intermediate Varchenko–Gelfand ideal
@@
For all $k$, $\sigma_{k}=\operatorname{dim}\operatorname{VG}_{k}(\mathcal{A},\mathbb{F})$. In particular, $\mathcal{A}$ satisfies Yoshinaga’s criterion if and only if $I_{2}=I_{r}$.
@@ 2.3 The Cordovil algebra
If $J_{2}=J_{r}$ (that is, if $\operatorname{C}(\mathcal{A},\mathbb{F})$ is quadratic), then $\mathcal{A}$ satisfies Yoshinaga’s criterion.
@@
From the sequence of surjections above, we see that the condition $J_{2}=J_{r}$ implies that $\bar{I}_{2}=\bar{I}_{r}$. Since $\operatorname{dim}\bar{I}=\operatorname{dim}I$ for any ideal $I\subset R$, this implies that $I_{2}=I_{r}$, which is equivalent to Yoshinaga’s criterion by Theorem 2.4. ∎
@@
The $D_{4}$ arrangement consists of the $12$ hyperplanes in ${\mathbb{R}}^{4}$ given by equations $x_{i}\pm x_{j}=0$ for $1\leq i<j\leq 4$. A Macaulay2 [GS] calculation easily shows that $I_{2}=I_{4}$, so $D_{4}$ satisfies Yoshinaga’s criterion. A similar calculation shows that the Cordovil ideal $J_{4}$ has minimal generators in degrees $2$ and $4$. That is, we have $J_{2}=J_{3}\subsetneq J_{4}=\bar{I}_{4}=\bar{I}_{3}=\bar{I}_{2}$.
@@ 3 Connections with other properties of arrangements
In this section, we prove that chordal  ⟹  Yoshinaga  ⟹  formal, and then collect known relationships between various properties of arrangements.
@@ 3.1 Chordality
We define a matroid to be chordal if, for every circuit $C$ of size at least $4$, there exist circuits $D_{1}$ and $D_{2}$ such that $\left|D_{1}\cap D_{2}\right|=1$ and
@@
The concept of chordality for graphs goes back to Berge [Ber69] and Dirac [Dir61]. Stanley noticed the connection between chordal graphs and supersolvability [Sta72, Example 2.7, Proposition 2.8]. Independently, Barhona and Grötschel introduced the notion of a chordal circuit as a way to characterize the facet-defining hyperplanes of the cycle polytope of a binary matroid [BG86, p.53]. Ziegler then showed that every binary supersolvable matroid not containing the Fano matroid is graphical [Zie91, Theorem 2.7]. Later Cordovil, Forge, and Klein showed that every binary supersovable matroid is chordal [CFK04, Theorem 2.2].
@@
If $\mathcal{A}$ is chordal, then $\mathcal{A}$ satisfies Yoshinaga’s criterion.
@@
We can check with Macaulay2 that $I_{2}=I_{3}$, hence Theorem 2.4 implies that $\mathcal{A}$ satisfies Yoshinaga’s criterion. The associated matroid are $20$ circuits, $5$ of which have three elements and $15$ of which have four elements. As there are only $10$ pairs of $3$-element circuits, $\mathcal{A}$ cannot be chordal.
@@ 3.2 Formality
If $\mathcal{A}$ satisfies Yoshinaga’s criterion, then $\mathcal{A}$ is formal.
@@
Suppose $\mathcal{A}$ is not formal. Then $V_{2}\supsetneq V$, so Lemma 3.5 tells us that $\mathcal{A}_{2}$ has more chambers than $\mathcal{A}$. Since chambers of $\mathcal{A}$ are in bijection with $\Sigma_{\operatorname{rk}\mathcal{A}}(\mathcal{A})$, this means that there exists a sign vector $\epsilon\in\Sigma_{\operatorname{rk}\mathcal{A}_{2}}(\mathcal{A}_{2})\setminus\Sigma_{\operatorname{rk}\mathcal{A}}(\mathcal{A})$. For every flat of rank 2, we have ϵ—_F ∈Σ_2((A_2)_F) = Σ_2(A_F), so $\epsilon\in\Sigma_{2}(\mathcal{A})$ by Proposition 3.4. But $\epsilon\not\in\Sigma_{\operatorname{rk}\mathcal{A}}(\mathcal{A})$, so $\mathcal{A}$ does not satisfy Yoshinaga’s criterion. ∎
@@
    $\mathcal{A}$ satisfies Yoshinaga’s criterion
@@ 3.3 Relationships
Yoshinaga’s criterion is closely-related to a number of other properties of arrangements. Below are several well-known arrangements, together with a summary of which properties they satisfy. Here OS refers to the Orlik–Solomon algebra, Cord refers to the Cordovil algebra, and both quad Cord and quad OS mean that the defining ideals of the correspinding rings are quadratically generated.
Arrangement $\mathbf{K(\pi,1)}$ free formal Yoshinaga quad Cordovil quad OS

### v2

# Cleanliness and the Varchenko–Gelfand algebra

@@
Abstract. A central question in the theory of hyperplane arrangements is when the complement of a complex arrangement is $K(\pi,1)$. Barkley and Speyer introduced a class of real arrangements that are called “clean,” and Yoshinaga proved that every real arrangement whose complexification is $K(\pi,1)$ is clean. We show that cleanliness is equivalent to a natural statement about the Varchenko–Gelfand ring, which in practice allows for fast calculation. We conclude with an investigation of the relationships between various properties of arrangements, including cleanliness and the $K(\pi,1)$ property.
@@ 1 Introduction
It is a long-standing open problem to determine which complex hyperplane arrangement complements are $K(\pi,1)$, meaning that their higher homotopy groups vanish. In the case where the hyperplane arrangement is the complexification of a real hyperplane arrangement, the homotopy type of the complement is determined by the associated oriented matroid [Sal87, Theorem 1], and therefore the $K(\pi,1)$ problem must have a combinatorial answer. See [FR00, FR87a, Yos24b] for surveys and partial results.
Recently, Yoshinaga proved that, if $\mathcal{A}$ is a real hyperplane arrangement and the complement of the complexification of $\mathcal{A}$ is $K(\pi,1)$, then $\mathcal{A}$ is clean in the sense of Barkley and Speyer [BS23, Yos24a]. In the first part of this paper, we give an algebraic reformulation of cleanliness, which we now describe.
Let $\mathcal{A}$ be a finite set of hyperplanes in a real vector space $V$, and let $\mathbb{F}$ be any field. The Varchenko–Gelfand algebra $\operatorname{VG}(\mathcal{A},\mathbb{F})$ is by definition the ring of locally constant $\mathbb{F}$-valued functions on the complement of the union of hyperplanes. This is a boring ring (it is isomorphic to a direct sum of one copy of $\mathbb{F}$ for each chamber), but it admits an interesting presentation whose generators are the Heaviside functions: there are two such functions for each hyperplane, taking the value 1 on one side of the hyperplane and 0 on the other side. The ring $\operatorname{VG}(\mathcal{A},\mathbb{F})$ is filtered, with the $p^{\text{th}}$ filtered piece consisting as functions that can be expressed as polynomials of degree at most $p$ in the Heaviside functions. The associated graded algebra, which is also called the Cordovil algebra, is isomorphic to the cohomology ring of the complement of the union of the subspaces $H\otimes{\mathbb{R}}^{3}\subset V\otimes{\mathbb{R}}^{3}$ [Mos17, DBPW24]. We say that the Varchenko–Gelfand algebra is quadratic if all relations among the Heaviside functions are generated by those of degree at most 2. Similarly, we say that the Cordovil algebra is quadratic if all relations among the corresponding generators are generated by those of degree 2. Our main results (Theorem 2.5 and Corollary 2.6) say that the following implications hold: $\operatorname{C}(\mathcal{A},\mathbb{F})$ is quadratic $\Longrightarrow$ $\operatorname{VG}(\mathcal{A},\mathbb{F})$ is quadratic $\Longleftrightarrow$ $\mathcal{A}$ is clean.
@@
At first sight, cleanliness (which is formulated combinatorially) might seem easier to work with than the condition that $\operatorname{VG}(\mathcal{A},\mathbb{F})$ is quadratic. In fact, our experience is that the algebraic condition is much faster to check, since computers are very good at using Gröbner bases to determine whether or not two ideals are equal. For example, let $\mathcal{A}$ be the arrangement whose normal vectors are given by the columns of the following matrix:
@@
It took about 1.65 seconds for Macaulay2 to determine that the Varchenko–Gelfand ideal is not quadratic. On the other hand, it took 3 days, 23 hours, 21 minutes, and 22 seconds for Sage to check cleanliness directly. [^4] Our Sage implementation could admit many improvements; but, even with considerable effort, it is unlikely that we could beat the time of the easy Macaulay2 calculation.
Our primary motivation for Theorem 2.5 is to be able to perform fast calculations, and in particular to probe the question of how close cleanliness is to the $K(\pi,1)$ property. The following examples illustrate the type of calculations that are made possible by our result.
@@
Let $\mathcal{A}$ be the arrangement in $\mathbb{R}^{6}$ consisting of all hyperplanes of the from $x_{i}=x_{j}$ for $1\leq i<j\leq 6$, together with the hyperplanes $x_{i}+x_{j}=0$ whenever jij-i is prime. (This is an intentionally unmotivated condition that is meant to produce a somewhat random arrangement lying in between the Coxeter arrangements of type $A_{5}$ and $D_{6}$.) The ring $\operatorname{VG}(\mathcal{A},{\mathbb{Q}})$ is not quadratic, and therefore $\mathcal{A}$ is not $K(\pi,1)$. This can be checked in Macaulay22 in about 30 seconds.
@@
When $t\in\{-1,0,1\}$, these arrangements have quadratic Varchenko–Gelfand algebras, and are therefore clean. Edelman–Reiner, however, show that these arrangements are not $K(\pi,1)$ [ER95, Theorem 2.1].
@@
This is the smallest known non-tame arrangement (see [Abe25] for background). Then $\operatorname{VG}(\mathcal{A};{\mathbb{Q}})$ is quadratic, thus $\mathcal{A}$ is clean. It is not known to the authors whether or not $\mathcal{A}$ is $K(\pi,1)$. Yoshinaga’s theorem gives supporting evidence that it could be.
Section 3 is devoted to relating cleanliness to other algebraic, topological, and combinatorial conditions. We define what it means for a matroid to be chordal, generalizing the notion of a chordal graph. We then say that $\mathcal{A}$ is chordal if its underlying matroid is chordal. We prove that every real, chordal arrangement is clean (Theorem 3.2). We also provide a proof (communicated to us by Paul Mücksch) that every clean arrangement is formal. The converses to these two theorems are false (Example 3.3 and Remark 3.8), but for graphical arrangements, chordality, formality, cleanliness, and the $K(\pi,1)$ property are all equivalent (Corollary 3.7). Finally, we provide a chart that illustrates the implications between various properties known to be related to the $K(\pi,1)$ property, including chordality, formality, cleanliness, and more.
@@ Acknowledgements
The authors would like to thank Nick Addington, Grant Barkley, Mike Falk, and Paul Mücksch for their valuable contributions.
2 Cleanliness and the Varchenko–Gelfand algebra
@@
2.1 Cleanliness
We begin by choosing coorientations of each element of $\mathcal{A}$. That is, for each $H\in\mathcal{A}$, we write $H^{+}$ to denote one of the two connected components of $V\setminus H$, and $H^{-}$ to denote the other one. To match the conventions in [BS23], we choose our coorientations in such a way so that the intersection of all of the positive half-spaces is nonempty. For any sign vector $\epsilon\in\{\pm\}^{\mathcal{A}}$ and any subset $S\subset\mathcal{A}$, let
@@ 2.1 Cleanliness
We say that $\mathcal{A}$ is clean if $\sigma_{2}=\sigma_{r}$.
Remark 2.1.
Our assumption that the intersection of all of the positive half spaces are nonempty implies that, if $\epsilon$ is a sign vector and $S\subset\mathcal{A}$ is a set of cardinality 3 with $H_{S}^{\epsilon}=\emptyset$, then the restriction of $\epsilon$ to $S$ either takes the value ++ twice and - once, or vice-versa. In the terminology of [BS23, Section 2.1], the sign vector $\epsilon$ is closed if there does not exist such an $S$ such that the restriction of $\epsilon$ to $S$ takes the value ++ twice, and it is coclosed if there does not exist such an $S$ such that the restriction of $\epsilon$ to $S$ takes the value - twice. The sign vector $\epsilon$ is biclosed if it is both closed and coclosed, which means that there is no set $S$ of cardinality 3 with $H_{S}^{\epsilon}=\emptyset$, or equivalently that $\epsilon\in\Sigma_{2}(\mathcal{A})$. Finally, $\epsilon$ is separable if $H_{\!\mathcal{A}}^{\epsilon}\neq\emptyset$, or equivalently if $\epsilon\in\Sigma_{r}(\mathcal{A})$. Thus cleanliness is precisely the statement that every biclosed sign vector is separable.
Our interest in clean arrangements comes from the following result [Yos24a, Theorem 5.1(2)].
@@
If $M_{2}(\mathcal{A})$ is $K(\pi,1)$, then $\mathcal{A}$ is clean.
@@ 2.2 The Varchenko–Gelfand algebra
In order to relate cleanliness to the Varchenko–Gelfand ring, we introduce a family of smaller ideals that sit inside the kernel of $\varphi$. For any $k$, we define the $k^{\text{th}}$ intermediate Varchenko–Gelfand ideal
@@
For all $k$, $\sigma_{k}=\operatorname{dim}\operatorname{VG}_{k}(\mathcal{A},\mathbb{F})$. In particular, $\mathcal{A}$ is clean if and only if $I_{2}=I_{r}$.
@@ 2.3 The Cordovil algebra
If $J_{2}=J_{r}$ (that is, if $\operatorname{C}(\mathcal{A},\mathbb{F})$ is quadratic), then $\mathcal{A}$ is clean.
@@
From the sequence of surjections above, we see that the condition $J_{2}=J_{r}$ implies that $\bar{I}_{2}=\bar{I}_{r}$. Since $\operatorname{dim}\bar{I}=\operatorname{dim}I$ for any ideal $I\subset R$, this implies that $I_{2}=I_{r}$, which is equivalent to cleanliness by Theorem 2.5. ∎
@@
The $D_{4}$ arrangement consists of the $12$ hyperplanes in ${\mathbb{R}}^{4}$ given by equations $x_{i}\pm x_{j}=0$ for $1\leq i<j\leq 4$. A Macaulay2 [GS] calculation easily shows that $I_{2}=I_{4}$, so $D_{4}$ is clean. A similar calculation shows that the Cordovil ideal $J_{4}$ has minimal generators in degrees $2$ and $4$. That is, we have $J_{2}=J_{3}\subsetneq J_{4}=\bar{I}_{4}=\bar{I}_{3}=\bar{I}_{2}$.
@@ 3 Connections with other properties of arrangements
In this section, we prove that chordal  ⟹  clean  ⟹  formal, and then collect known relationships between various properties of arrangements.
@@ 3.1 Chordality
We define a matroid to be chordal if, for every circuit $C$ of size at least $4$, there exist circuits $D_{1}$ and $D_{2}$ such that $\left|D_{1}\right|,\left|D_{2}\right|\geq 3$, $\left|D_{1}\cap D_{2}\right|=1$, and
@@
The concept of chordality for graphs goes back to Berge [Ber69] and Dirac [Dir61]. Stanley noticed the connection between chordal graphs and supersolvability [Sta72, Example 2.7, Proposition 2.8]. Independently, Barhona and Grötschel introduced the notion of a chordal circuit as a way to characterize the facet-defining hyperplanes of the cycle polytope of a binary matroid [BG86, p.53]. Ziegler then showed that every binary supersolvable matroid not containing the Fano matroid is graphical [Zie91, Theorem 2.7]. Later Cordovil, Forge, and Klein showed that every binary supersolvable matroid is chordal [CFK04, Theorem 2.2].
@@
If $\mathcal{A}$ is chordal, then $\mathcal{A}$ is clean.
@@
We can check with Macaulay2 that $I_{2}=I_{3}$, hence Theorem 2.5 implies that $\mathcal{A}$ is clean. The associated matroid are $20$ circuits, $5$ of which have three elements and $15$ of which have four elements. As there are only $10$ pairs of $3$-element circuits, $\mathcal{A}$ cannot be chordal.
@@ 3.2 Formality
If $\mathcal{A}$ is clean, then $\mathcal{A}$ is formal.
@@
Suppose $\mathcal{A}$ is not formal. Then $V_{2}\supsetneq V$, so Lemma 3.5 tells us that $\mathcal{A}_{2}$ has more chambers than $\mathcal{A}$. Since chambers of $\mathcal{A}$ are in bijection with $\Sigma_{\operatorname{rk}\mathcal{A}}(\mathcal{A})$, this means that there exists a sign vector $\epsilon\in\Sigma_{\operatorname{rk}\mathcal{A}_{2}}(\mathcal{A}_{2})\setminus\Sigma_{\operatorname{rk}\mathcal{A}}(\mathcal{A})$. For every flat of rank 2, we have ϵ—_F ∈Σ_2((A_2)_F) = Σ_2(A_F), so $\epsilon\in\Sigma_{2}(\mathcal{A})$ by Proposition 3.4. But $\epsilon\not\in\Sigma_{\operatorname{rk}\mathcal{A}}(\mathcal{A})$, so $\mathcal{A}$ is not clean. ∎
@@
    $\mathcal{A}$ is clean
@@ 3.3 Relationships
Below are several well-known arrangements, together with a summary of which properties they satisfy. Here OS refers to the Orlik–Solomon algebra, Cord refers to the Cordovil algebra, and both quad Cord and quad OS mean that the defining ideals of the corresponding rings are quadratically generated.
Arrangement $\mathbf{K(\pi,1)}$ free formal clean quad Cordovil quad OS